A particle with mass 2.47 kg oscillates horizontal
Answer #1

n = number of cycles = 71

t = duration = 121 s

f = frequency = n/t = 71/121 = 0.587 hz

w = angular frequency = 2\pif = 2 (3.14) (0.587) = 3.7 rad/s

A = amplitude = 0.913 m

Vmax = Aw = 0.913 x 3.7 = 3.4 m/s

m = 2.47 kg

w = sqrt(k/m)

3.7 = sqrt(k/2.47)

k = 33.81 N/m

Umax = (0.5) k A2 = (0.5) (33.81) (0.913)2 = 14.1 J